### v1

# Modeling the Carbon Footprint of HPC: The Top 500 and EasyC

Abstract
Climate change is a critical concern for HPC systems, but GHG protocol carbon-emission accounting methodologies are difficult for a single system, and effectively infeasible for a collection of systems. As a result, there is no HPC-wide carbon reporting, and even the largest HPC sites do not do GHG protocol reporting.
We assess the carbon footprint of HPC, focusing on the Top 500 systems. [^1] This work has been accepted at the ACM/IEEE Workshops of the International Conference for High Performance Computing, Networking, Storage and Analysis (SC Workshops ’25), November 16–21, 2025, St Louis, MO, USA. The key challenge lies in modeling the carbon footprint with limited data availability.
With the disclosed Top500.org data, and using a new tool, EasyC, we were able to model the operational carbon of 391 HPC systems and the embodied carbon of 283 HPC systems. We further show how this coverage can be enhanced by exploiting additional public information. With improved coverage, then interpolation is used to produce the first carbon footprint estimates of the Top 500 HPC systems. They are 1,393.7 million MT CO2e operational carbon (1 Year) and 1,881.8 million MT CO2e embodied carbon. We also project how the Top 500’s carbon footprint will increase through 2030.
A key enabler is the EasyC tool which models carbon footprint with only a few data metrics. We explore availability of data and enhancement, showing that coverage can be increased to 98% of Top 500 systems for operational and 80.8% of the systems for embodied emissions.
Index Terms:
Top500, HPC, Sustainability, Carbon Emissions
I Introduction
Global warming is accelerating at an alarming rate[1], and computing’s increased use is an important and growing contributor. Data center global power demand is predicted to increase 165% by 2030 compared to 2023[2]. A significant part of this growth is driven by emerging compute and hardware-intensive applications, such as large language models [3, 4, 5] and the increasing use of scientific simulations[6]. While these applications offer substantial potential for scientific progress, they also carry a considerable carbon footprint. Within the High Performance Computing (HPC) community, there is rising concern as evidenced by increased scientific meetings, and a growing body of literature [7, 8, 9, 10, 11, 12]. This interest is driven both by the civic/environmental awareness of the HPC community, and more starkly by the rapid growth of the carbon footprint of HPC systems which are consuming more energy and producing higher carbon emissions due to continued expansion [13, 14]. This concern is broad-based, despite little broad analysis of HPC carbon footprint.
Despite this, carbon emission reporting for HPC systems is rare. For example, none of the Top 25 computing systems (Top 500 list) report their carbon footprint. GHG protocol carbon emission accounting methodologies [15] require far too much effort, and more problematically a broad collection of accurate data inputs. Despite that, GHG protocol approaches are used by many large companies, including tech companies[16, 17, 18, 19], at large cost and with outcomes of questionable value.
Academic research studies have skirted the limits of significant per-system effort and extensive data requirements (under the GHG protocols), compromising either the rigor of studies, guessing or estimating many of the input data metrics [11, 10, 20]. External estimates of computer systems data metrics are even more difficult, making objective academic study challenging and often inaccurate. So, these existing examples provide no way to report the carbon footprint of a specific identifiable set of systems, such as the Top500, government-funded research computing facilities, or scientific HPC platforms in Europe. This is unfortunate because to reduce emissions from HPC systems, we need widespread carbon footprint assessments. Achieving this at scale requires methods that work with limited data and low human effort.
We use the Top500 November 2024 list[14] as a proxy to evaluate the carbon emissions of HPC systems. We perform a broad analysis of the Top 500 systems, enabled by the EasyC[21] tool that dramatically reduces the detailed data requirements and effort, compared to the GHG protocol. The results characterize the operational and embodied carbon for this collection of systems, representative of the HPC computing community. We further examine multiple data input scenarios, assessing coverage and the change that new data can make to carbon footprint models (sensitivity). Finally, we use the data to project the future changes in HPC carbon footprint, and important metrics such as compute-to-carbon footprint ratio.
Specific contributions include:
  • The first assessment of Top500 carbon footprint. Annual operational carbon is 1,393.7 million MT CO2e (equivalent to 325k gasoline-powered vehicles annual emissions). Embodied carbon is 1,881.8 million MT CO2e (equivalent to 439k gasoline-powered vehicles annual emissions).
  • An assessment of a carbon modeling tool, EasyC, that shows the carbon footprint (operational and embodied) of 56.6% of the Top 500 systems can be captured using only the data from Top500.org.
  • With additional public information, the coverage increases to 98% for operational carbon and 80.8% for embodied carbon. The carbon footprint of the remaining systems can be further computed through interpolation. Thus the carbon footprint of the entire Top500 can be computed with Top500.org data, other publicly available sources, and interpolation.
  • Projection of the Top 500 carbon footprint thru 2030. Operational carbon footprint continues to increase rapidly (10.3%/year), reaching 1.8x its 2024 level by 2030. Embodied carbon grows slower (2%/year), reaching 1.1x.
II Problem
Carbon emissions from HPC systems are increasing, but sustainability reporting for these systems is basically non-existent. Existing carbon accounting methodologies require substantial data and human effort, making them impractical for widespread use. Research computing facilities often operate under tight budgets, with staff juggling multiple roles to support a diverse and demanding user base. As a result, they have limited capacity to take on additional sustainability efforts.
We focus on the challenge of assessing the carbon footprint of scientific/HPC systems. How can it be done comprehensively? How can we create a basis for regular reporting? A solution for it must have the following attributes:
  • Achieve broad coverage, including the full range of HPC systems.
  • Employ only limited data (a few data metrics) that are readily available
  • Require only modest human effort, as many HPC organizations are short-staffed
  • Provide a ”gentle slope”, allowing the addition of data metrics for unusual circumstances.
And we believe that, to be practicable, carbon footprint reporting for each system should require less than a person-hour of effort per year.
III Approach
We assess the carbon footprint of HPC systems using the Top500 November 2024 list [14] as a representative set of systems. [^2] We expect to update for the June 2025 Top 500 list by the time of the workshop. Our approach builds on an existing tool for carbon-footprint modeling for computer systems called EasyC. It requires modest effort, low data, and provides high quality footprint reporting [21], see Figure 1.
We begin with data from top500.org; however, this information is incomplete for many systems. As many as 300-500 systems are missing specific pieces of required Top 500 reporting information (see Figure 2). To improve coverage and global applicability, we augment it with other publicly available sources (e.g., web page scraping) and refine the framework accordingly. Using these datasets, we conduct sensitivity analyses and assess the carbon footprint of the systems, as well as project future changes in the HPC carbon footprint.
Our results show that EasyC substantially increases the carbon-assessment coverage of these systems, reduces the data collection effort, and supports customization for diverse system configurations.
IV Evaluating the Carbon Footprint of HPC
In the following sections, we assess the carbon emissions of HPC systems using the Top500 list as a proxy. First, we apply a powerful carbon footprint modeling tool (EasyC) to the available data from top500.org, and show how significant additional work is required to construct acceptable coverage of the 500 systems. Second, we use the improved coverage to compute a high-quality assessment of the Top 500’s carbon footprint. Finally, we project the carbon footprint of the Top 500 for 2025-2030.
IV-A Carbon Footprint of the Top500 HPC Systems
We assess the carbon footprint of the Top 500 supercomputing systems in the world. To our knowledge, this has never been done before. Thus, this study represents the first estimate of the global footprint of high-performance computing’s carbon footprint.
We use data from the Top 500 website. As part of its performance ranking, top500.org collects both performance and a variety of structural features of the systems (processor model, co-processor model, total cores, etc.) as well as several performance attributes ( Rmax[TFlop/s], Rpeak[TFlop/s], Nmax, and others). The structural and performance information is ”high quality” for all 500 systems, because its a requirement for inclusion in the Top 500. However there are data gaps that cause problems with carbon-footprint assessment (see Figure 2).
Refer to caption EasyC: The EasyC tool calculates the carbon footprint (operational and embodied), using sophisticated models that identify the key data metrics required. It needs just 7 key data metrics[21]. This differs from the widely used GHG Protocol that can require hundreds of metrics.
Figure 1: EasyC Carbon Footprint Tool
Figure 2: Structural information reported for different data items in the Top 500. Note: ’None’ indicates that all information is reported.
Based on the available Top 500 structural information, we use the EasyC tool to assess the carbon footprint for all the systems that sufficient data is available. This is 391/500 systems for operational carbon and 283/500 for embodied carbon. Figures 3(a) and 3(b) shows operational (1 Year) and embodied carbon versus rank. This coverage, while the best to date, is still too poor to construct an overall Top 500 assessment because it covers little more than half of the systems. Further, the systems present and missing are highly skewed; the highest-performing systems, the largest contributors, are disproportionately missing.
(a) Operational Carbon
(b) Embodied Carbon
Figure 3: Top 500 Carbon Footprint vs Rank (Top500.org data)
Steps to Increase Coverage How to increase coverage? Interpolation is not robust with such a sparse collection. We turn to the core idea of EasyC, which is that intelligent use of just a few key data metrics can be used to create accurate carbon modeling (see Figure 1). It’s natural to ask how and what additional sources of data could be exploited to enable better coverage. Specifically, we consider two data metric scenarios: (1) Baseline, uses only data available on Top500.org, and (2) Baseline+Public Info, supplements the data with publicly available information on other web sites. We first assess coverage, and then quantitative carbon metrics.
TABLE I: Data EasyC requires, but not available on Top500.org, public disclosure
Type # Systems Incomplete # Systems Incomplete
[Top500.org] [Other Public]
Operation Year 0 0
# of Compute Nodes 209 86
# of GPUs 209 86
# of CPUs 0 0
Memory Capacity 499 292
Memory Type 500 292
SSD Capacity 500 450
System Util (opt.) 500 497
Annual Power Consumed (opt.) 500 492
# Data Metrics Missing 9 9
Because of their leading edge nature, the Top 500 systems are particularly challenging to analyze. Even the minimum data required to estimate carbon emissions is missing for many systems. Table I shows the extent of data incompleteness when considering both top500.org and other publicly available sources. To compute operational carbon requires actual power consumption and power source. Embodied carbon requires system configuration and component information. Exacerbating this, some systems use early or unique compute devices (eg MI300A, Fugaku A64FX, Sunway SW26010, etc.) particularly for embodied carbon. Essentially all of the Top 500 do not report actual power consumption. Many of the systems are commercially operated or government systems with no public information available. Only a few are ”open science” systems with good configuration information available.
Further, we found that none of the systems provided reporting under the GHG protocol (so there are no reports to compare with).
(a) Operational Carbon
(b) Embodied Carbon
Figure 4: Carbon Footprint Reporting Coverage
In Figure 4, we show how the coverage of systems increases varies with different data metric availability. Using the GHG detailed carbon accounting method, few of the Top 500 systems reports operational and NONE report embodied. As we add public data, coverage of the Top 500 operational carbon increases from 78% of systems on top500.org using only the data available there. Including publicly available data (number of compute nodes, number of GPUs, number of CPUs, memory capacity, and SSD capacity), this coverage jumps to 98%, allowing us to estimate almost all systems’ operational emissions. For embodied carbon emissions, we see a 1.43x improvement in coverage for estimating embodied carbon emissions. This is because data crucial for embodied carbon estimates like the number of nodes, accelerators, or processors is often missing from top500.org.
(a) With Top500.org data
(b) With Top500.org and other public data
Figure 5: Operational carbon coverage of EasyC for varied system ranks (two data input scenarios).
(a) With Top500.org data
(b) With Top500.org and Public data
Figure 6: Embodied carbon coverage of EasyC for varied system ranks (two data input scenarios).
Figures 5 and 6 present the varying coverage across Top500 system ranking ranges. The operational carbon emissions of many Top 500 systems can be estimated using data from top500.org. Significant gaps emerge surprisingly high in the rankings 26-50, 51-75, and 76-100. The main reason is that many systems do not report their power consumption, which is essential when information on the number of compute nodes and GPU nodes is unavailable. The addition of public information fills these gaps, rendering nearly full coverage.
For embodied carbon, the situation is much worse. For many systems in the Top 150, there was insufficient data to make an estimate, even with the limited data requirements of EasyC. The major reason for this is that top systems today make heavy use of an increasingly diverse set of accelerators (e.g., Nvidia, AMD, many versions). Top500.org does not capture adequate accelerator information. For most of the systems 151-500 systems, they are CPU-based. Because they do not have accelerators, the number of CPU cores per node and total CPU cores that are captured at top500.org are sufficient for embodied carbon. We can see in Figure 6b that incorporating additional publicly available data on which accelerators were used is essential to improve coverage.
IV-B Constructing an Estimate of the Full Top 500
With much higher coverage, we employ simple statistical techniques to construct an assessment of carbon footprint for the full Top 500 supercomputers.
To fill the gaps of 100’s of systems, we interpolate the carbon footprint for the systems missing data using the average of the nearest 10 peers (5 lower and 5 higher) in the Top 500.333If the peers are also incomplete, we use the next closest peers. The data in Figure 7a represent the carbon footprint of HPC. On the left, we present the EasyC carbon footprint for all the systems covered by Top500.org plus other public information. On the right, we include those and add the interpolated systems. Continuing, we also compute the average operational and embodied carbon of systems in the systems with enough data (for EasyC) and for the full 500, completed by interpolation. This brings the Operational carbon for the Top 500 to 1,393.7 million MT CO2e, adding the missing 10 systems increased operational footprint by only 1.74%.
(a) Total Carbon Footprint
(b) Average System Carbon Footprint
Figure 7: Total Operational carbon (1 year) and Embodied Carbon for systems covered by Top500 + other public data and then full Top 500, interpolating for the missing systems. Then average carbon for the same sets of systems.
Adding the missing 96 systems increased Embodied carbon to 1,881.8 million MT CO2e, an increase of 23.18%. This operational carbon is equal to one year’s emissions for 325,000 gasoline-powered vehicles or 3.5 billion vehicle miles. The embodied carbon is equivalent to one year’s emissions for 439,000 gasoline-powered vehicles or 4.8 billion passenger miles. Of course the embodied carbon is 1-time for the lifetime of the computer system so it would be smaller if annualized. In Figure 8b, the average operational and embodied carbon for Top 500 systems is presented. Each is millions of MT CO2e, comparable to that of thousands of homes.
In Figure 7, we report the aggregate carbon footprint for the Top 500, both with the systems for which we have sufficient data. For 490 systems, the total operational carbon is 1,369.9 Million MT CO2e, and for 404 systems the total embodied carbon is 1,527.7 Million MT CO2e.
(a) Operational Carbon
(b) Embodied Carbon
Figure 8: Top 500 Systems Carbon Footprint, by rank. Full assessment, includes data generated by EasyC on top500.org and other public web data, augmented by interpolation.
Figure 9: How Operational and Embodied carbon footprint changes between Baseline (Top 500 data) and Baseline+PublicInfo.
The complete Top 500 results for each system are shown in Figure 8. It includes all of the data generated by EasyC on top500.org and public web data, augmented by 10 systems for operational and 96 systems for embodied modeled by interpolation. In short, it includes a model for all 500 systems in the Top 500 list.
As a sensitivity study, we show the difference that adding public information made in Figure 9. For operational carbon, we can see that adding public information does not change the overall results significantly. For operational carbon, we see refinement to a narrower region average carbon intensity (ACI) of power, which can increase or decrease by as much as 77.5%. But even a number of these do create a large change, the total change for the entire Top 500 is only 2.85% (38 million MTCO2e). For embodied carbon, there are larger changes, mostly increasing the carbon footprint, but these have small contribution. The biggest change is due to large systems where no estimate was previously possible (not shown). Overall this produces and increase of 670.48 million MTCO2e, for an 78% change overall. The use of novel accelerators, not documented in public information, is the largest problem. Approximating these accelerators with mainstream GPUs produces systematic underestimates of silicon size.
Overall, our results show that operational carbon emissions can typically be derived directly from top500.org. This is true across the range from 1-500, and for systems with and without accelerators. Embodied carbon is more difficult, as systems with accelerators are difficult due to their diversity, and growing number of types. But with additional other public data, our results show that EasyC can compute carbon emissions of almost all the systems on Top500, and with little effort.
IV-C How Carbon Footprint will Change for the Top 500?
(a) Operational Carbon
(b) Embodied Carbon
Figure 10: Projected Top 500 Operational and Embodied Carbon (2025-2030)
We project the carbon footprint of the Top 500 HPC systems for 5 years. An average of 48 systems was added to each new list in each cycle, over the past two years. With this turnover comes a 5% increase in operational carbon, and 1% increase in embodied. Annualized, this is 10.3% growth in operational and 2% growth in embodied carbon.
We project carbon footprint of the Top 500 based on these growth rates (see Figure 10) Operational carbon’s 10.3% annual growth reflects both increasing computing power represented in the Top 500 systems, but also the slowing benefits of Moore’s Law and the reality of post-Dennard scaling, as power consumption continues to increase rapidly. By 2030, Top 500’s operational carbon is nearly double that of 2024. Clearly architectural customization and accelerators is not enough to reduce power consumption.
In comparison, the embodied carbon emissions is growing slower (only 1.02x or 2% per year), suggesting its a lesser concern. However, we would advocate caution as embodied carbon is heavily influenced by storage system – not a focus of this study.
(a) Operational Carbon
(b) Embodied Carbon
Figure 11: Projected Top 500 Operational and Embodied Performance to Carbon Ratio (2025-2030)
Compute scaling has benefitted from dramatic increases on performance per unit (size, power, etc.) over more than seven decades [22, 23]. In recent times, those microelectronics scaling drivers have slowed or even stopped. With a sustainability focus, we consider performance per carbon footprint for the Top 500 HPC systems (see Figure 11).
The projection shows a continued increase in perf/carbon, but at an annual rate of 0.2 PFlop/s million MT CO2e. While there is still progress, this is dramatically slower than what we experienced for decades under Dennard scaling, and parallelism scaling where we’d expect 2x performance / unit power every 18-month generation (plotted for comparison). Note that the current increase in performance / unit carbon is not sufficient to compensate for the rapid growth of use of computing. As the net growth of total carbon footprint of the Top 500 continues at 10.3%/annum (see Figure 10).
V Discussion and Related Work
V-A Organizational Reporting (GHG Protocol)
Many organizations including Google, Microsoft, TSMC, Intel, United Airlines, Emirates, Harvard University, Argonne National Laboratory, and others publish annual sustainability reports[17, 18, 24, 25, 26, 27, 28, 29] based on the GHG Protocol[15]. These reports take a broad view, covering carbon emissions from construction, employee commuting, purchased goods and services, capital goods, IT, and more. However, their scope is generally at the organizational level rather than the level of individual computing systems, making it difficult to directly assess the carbon footprint of specific computing systems which is the focus of this work.
V-B Studies/Reporting of Individual Computer Systems
Several recent academic studies have estimated the carbon emissions of specific systems[30, 10, 12] However, their methodologies are based on the GHG Protocol, which requires extensive data inputs and significant effort for each system. Thus, their approaches cannot be easily adapted to assessing a wide range of computing systems, the focus of this paper.
V-C GHG Protocol vs EasyC
The GHG Protocol [15] is an accepted standard for carbon accounting. It is based on comprehensive data collection and tabulation. This approach means that accuracy is dependent on every piece of data needed, a classical accounting approach. Further, exhaustive data collection and tabulation is a major effort, and invites the include of inaccurate data.
Research computing systems are complex aggregations of thousands of devices, each comprised of hundreds of components, produced by complex supply chains. As a result, each device and component can have important differences in life-cycle history and performance (e.g. power efficiency). For example, a system may use DRAMs manufactured in South Korea, Japan, and the United States. Since each manufacturing site differs, even DRAM with the same capacity can vary in components and materials used. Further, even if possible to obtain data for each component and device, each inclusion incorporates additional data inaccuracies into the larger computation. There’s no reason to believe that ”statistical variation” will reduce error, as these inputs are rife with systematic error. For example, errors can arise from using outdated or inaccurate carbon intensity data for electricity, often due to limited public availability, inconsistent time granularity, or omission of on-site energy sources.
Finally, the effort required to do a diligent GHG Protocol computation for a computer system is daunting, perhaps requires weeks of effort, and thus may be impractical for many research computing facilities that are often staffing-limited.
However, EasyC uses a statistical approach to identify the key metrics which are predictive of carbon footprint. As a result, EasyC reduces the data metrics needed to only seven, dramatically lowering effort required. This enables internal assessment of research computing systems, and also external assessment of groups of computing systems such as the Top 500.
VI Summary and Future Work
Detailed carbon emission computation methodologies fail to achieve high accuracy even with comprehensive data and significant effort. This makes it impossible for research computing facilities with limited resources to use it. However, our evaluation on Top500 systems is a first large-scale assessment of HPC’s carbon footprint. The fact that it was even feasible demonstrates EasyC’s ability to compute carbon emissions for diverse heterogeneous systems, unlike more detailed approaches based on GHG protocol and life-cycle assessment.
Going forward, we hope that this work will lead to widespread carbon emission reporting of computing systems. For example, we would like to model carbon footprint for all of the US National Science Foundation ACCESS scientific computing sites, those of the US Department of Energy, or of similar such systems in Europe or China.
Acknowledgment
We thank the anonymous reviewers for the insightful reviews. This work is supported in part by NSF Grants OAC-2442555, CNS-2325956, and the VMware University Research Fund. We also thank the Large-scale Sustainable Systems Group members for their support of this work!